complete ordered pair solutions

3x+4y+12 (0,),(,3\4) ,(,0),(8\3,)

Answer by checkley71(8405)

3X+4Y+12=0
3*0+4Y+12=0 OR 4Y=-12 OR Y=-12/4 OR Y=-3 THUS(0,-3)IS THE ANSWER.
3X+4(3/4)+12=0 OR 3X+3+12=0 OR 3X=-15 OR X=-15/3 OR X=-5 THUS(-5,3/4) IS THE ANSWER.
3X+480+12=0 OR 3X+12=0 OR 3X=-12 OR X=-12/3 OR X=-4 THUS(-4,0) IS THE ANSWER.
4*8/3+4Y+12=0 OR 32/3+4Y+12=0 OR 4Y=-12-32/3 OR 4Y=(-36-32)/3 OR 4Y=-68/3 OR
Y=-68/3*1/4 OR Y=-68/12 OR Y=-17/3 THUS(8/3,-17/3) IS THE ANSWER.